40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 133 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  10284   Tue Jul 29 07:54:34 2014 SteveUpdateSUSETMX damping restored

ETMX sus damping restored

  10283   Mon Jul 28 17:53:00 2014 AkhilUpdateGeneralWork plan for the Upcoming weeks- FOL Project

 [Akhil, Harry]

Work Completed :

 Frequency Counter:

  • Interfacing with the Raspberry Pi
  • Characterization of the FC:          

                                   - Transfer Function 

                                   - Quantization Noise Estimation

Temperature Actuator:

  • Measurement of the Transfer Function

EPICS and Channel Readout:

  •  Creating a new Channel Access Server(SoftIOC)
  •   Piping data from FC into created channels.              

Frequency Offset Locking(FOL) Box Design and Plan:

  • Planning and selection of place for installation.
  • Preparation of the box and routing cables.

                

Work Plan for Upcoming Weeks:

  • Calibration of the Thermal Actuator TF and PID loop design.
  • Channel Testing after installation of the FOL box inside the 40m.
  • Optics:
    • Measure beam profiles of AUX lasers and PSL.
    • Design coupling telescope, given space constraints at end tables
    • Couple lasers into fibers
    • Connect fibers from lasers to fiber coupled Beam Combiner and Photodiode.
  • Testing of FOL loop after installation of the complete system.

 

 

  10282   Mon Jul 28 17:25:32 2014 HarryUpdateGeneralFiber Mode With Collimators

 Purpose

We want a measurement of the fiber modes at either end, with the collimators, because these will be the modes that we'll be trying to match in order to couple light into the fibers, for FOL and/or future projects.

Measurement

In order to measure these modes, I used the beam profiler (Thorlabs BP 209-VIS) to take measurements of the beam diameter (cut off at 13.5% of the amplitude) along the optical axis, for each of the fiber ends.

The ends are arbitrarily labelled End 1 and End 2.

For each measurement, the fibers were coupled to roughly 30%, or 25mW at the output.

Regarding the issue of free rotation in the collimator stages: while End 1 was relatively stable, End 2 tended to move away from its optimal coupling position. In order to correct for this, I chose a position where coupling was good, and repositioned the stage to that coordinate (124 degrees) before taking each measurement.

The data were then entered into A La Mode, which gave waist measurements as follows:

End 1--- X Waist: 197um at Z = 4.8mm       Y Waist: 190um at Z = 13.6mm

End 2--- X Waist: 192um at Z = 7.4mm       Y Waist: 190um at Z = 6.0mm

end1Profiles.pngend2Profiles.png

A La Mode code is attached in .zip file

Moving Forward

These are the types of profiles that we will hopefully be matching the PSL and AUX lasers to, for use in frequency offset locking.

More characterization of the fibers is to follow, including Polarization Extinction Ratio.

We also hope to be testing the overall setup soon.

 

 

Attachment 3: FiberModeWCollimators.zip
  10281   Mon Jul 28 16:34:02 2014 AkhilUpdateGeneralCalibration of measured Thermal Actuator TFs

 To calibrate the measured TFs and characterize the thermal actuator for the FOL loop, we [ Me, Eric Q, Koji ] made the TF measurements of PZT response by giving a  disturbance to the position of  each of X and Y arm ETM  and ITM.

In order to make reasonable conclusions, the measurements were done at frequencies greater than 20 Hz (assuming the PZT response to be flat till a few KHz), which is out of the  bandwidth of the control loops operating for other noises at low frequencies, so that we can get the response only( mainly) due to the disturbance of the masses. 

 For this measurement , a Sine sweep excitation was given as an input to one of the test mass and PZT actuation signal was monitored. The channels used for the measurement are: 

Input( Mirror displacement):

ITMX- C1:SUS-ITMX_LSC_EXC

ETMX- C1:SUS-ETMX_LSC_EXC

ITMY- C1:SUS-ITMY_LSC_EXC

ETMY- C1:SUS-ITMX_LSC_EXC

Output ( PZT Response):

C1:ALS-Y_SLOW_SERV_IN1

The units of the TF of these measurements are not calibrated  and are in count/count. For this I will use the ITMX and ITMY calibration values from Izumi's Elog. I will also make some calculations and post in the calibrations of ETMX and ETMY in a separate elog.

I am now estimating the calibrated Thermal Actuator TF and will estimate the location of poles and zeroes to build the PID loop. I will elog the final calibrated TFs in my next elog.

The attached are the Bode Plots  for ETM and ITM for X and Y arms.

Attachment 1: mirrorTF2.pdf
mirrorTF2.pdf
  10280   Mon Jul 28 10:42:43 2014 NichinUpdateElectronicsDemodulator board's characterization

 I used vector fitting to fit the transfer functions between RF input and PD RF MON of demodulator boards. These fittings can certainly do a lot better on LISO, but for the time being I will assume these to be good enough and change the main PDFR scripts to calibrate out this factor and get a decent reading of PD transimpedance. Then it will just be a matter of changing the transfer function parameters. A lot of work needs to be done on the PDFR interface and plot features.

Attached: The plots showing data and fits.

Attachment 1: Demod_Fit.pdf
Demod_Fit.pdf Demod_Fit.pdf Demod_Fit.pdf Demod_Fit.pdf Demod_Fit.pdf Demod_Fit.pdf Demod_Fit.pdf Demod_Fit.pdf
  10279   Sat Jul 26 15:30:15 2014 Joseph AreedaUpdateComputer Scripts / ProgramsNDS2 server propem on megatron

The NDS2 server on megatron was unresponsive for what i think was the last couple of days.

The NDS the log file (~nds2mgr/logs/nds2-201407151045.log) started reporting "Stage: parser output queue is full." at 2014.7.24 14:47:54 also there are 16 connections still not closed with LindmeierLaptop.cacr.caltech.edu (131.215.146.102) with 15 of them in CLOSE_WAIT. 

To identify these zombie sockets we use "netstat -an | grep 31200"

The server was in a condition that /etc/init.d/nds2 stop didn't work and the process had to be manually kill -9'ed and then about 3 or 4 minutes later the zombie sockets were gone at /etc/init.d/nds2 start was used to restart the server.

The LindemejerLaptop was using pynds to get a bunch of channels at once to test drive a streaming visualization code for glitches.  It's unclear whether this bumped into a server limitation.  We have seen similar states in ldvw that seem to be the result of errors which result in client-server connections not being closed properly, leaving data in an output buffer causing Linux to wait for the other side to empty the buffer.

  10278   Sat Jul 26 14:45:33 2014 GabrieleMetaphysicsASCResponse of POP QPD

 Koji asked me to perform a simulation of the response of POP QPD DC signal to mirror motions, as a function of the CARM offset. Later than promised, here are the first round of results.

I simulated a double cavity, and the PRC is folded with parameters close to the 40m configuration. POP is extracted in transmission of PR2 (1ppm, forward beam). For the moment I just placed the QPD one meter from PR2, if needed we can adjust the Gouy phase. There are two QPDs in the simulation: one senses all the field coming out in POP, the other one is filtered to sense only the contribution from the carrier field. The difference can be used to compute what a POP_2F_QPD would sense. All mirrors are moved at 1 Hz and the QPD signals are simulated:

pop_qpd_all.png

This shows the signal on the POP QPD when all fields (carrier and 55 MHz sidebands) are sensed. This is what a real DC QPD will see. As expected at low offset ETM is dominant, while at large offset the PRC mirrors are dominant. It's interesting to note that for any mirror, there is one offset where the signal disappears.

pop_qpd_carrier.png

This is the contribution coming only from the carrier. This is what an ideal QPD with an optical low pass will sense. The contribution from the carrier increases with decreasing offset, as expected since there is more power.

pop_qpd_sb.png

Finally, this is what a 2F QPD will sense. The contribution is always dominated by the PRC mirrors, and the ETM is negligible.

The zeros in the real QPD signal is clearly coming from a  cancellation of the contributions from carrier and sidebands.

The code is attached.

Attachment 4: foldeddoublecavity.mist
classname FoldedDoubleCavity

# parameters
const Pin  1                # input power
const Lprc 6.752            # power recycling cavity length
const d_BS_PR3 0.401        # folding mirror distances
const d_PR2_PR3 2.081
const d_PRM_PR2 1.876
const c 299792458           # speed of light
const fmod 5*c/(4*Lprc)     # modulation frequency, matched to Lprc
... 51 more lines ...
Attachment 5: pop_qpd.m
% compile simulation class
clear classes
m = MIST('foldeddoublecavity.mist');

% create simulation object
s = FoldedDoubleCavity(8);

% set angulat motion
s.PRM.setMotionShape('pitch');
s.PR2.setMotionShape('pitch');
... 85 more lines ...
  10277   Sat Jul 26 14:35:28 2014 AkhilUpdateGeneralData Acquisition from FC into EPICS Channels

Quote:

Quote:

 I succeeded in creating a new channel access server hosted on domenica ( R Pi) for continuous data acquisition from the FC into  accessible channels. For this I have written a ctypes interface between EPICS and the C interface code to write data into the channels. The channels which I created are:

C1:ALS-X-BEAT-NOTE-FREQ

C1:ALS-Y-BEAT-NOTE-FREQ

 

The scripts I have written for this can be found in:

db script in:     /users/akhil/fcreadoutIoc/fcreadoutApp/Db/fcreadout.db

 Python code:  /users/akhil/fcreadoutIoc/pycall

C code:          /users/akhil/fcreadoutIoc/FCinterfaceCcode.c

I will give the standard channel names(similar to the names on the channel root)once the testing is completed and confirm that data from FC is consistent with the C code readout. Once ready I will run the code forever so that both the server and data acquisition are in process always.

Yesterday, when I set out to test the channel, I faced few serious issues in booting the raspberry pi. However, I have backed up the files on the Pi and will try to debug the issue very soon( I will test with Eric Q's R Pi).

To run these codes one must be root ( sudo python pycall, sudo ./FCinterfaceCcode)  because the HID- devices can be written to only by the root(should look into solving this issue). 

Instructions for Installation of EPICS, and how to create channel server on Pi will be described in detail in 40m Wiki ( FOLL page).

 

controls@rossa|~ 2> ls /users/akhil/fcreadoutIoc
ls: cannot access /users/akhil/fcreadoutIoc: No such file or directory
controls@rossa|~ 2> 

This code should be in the 40m SVN somewhere, not just stored on the RPi.

I'm still confused why python is in the mix here at all.  It doesn't make any sense at all that a C program (EPICS IOC) would be calling out to a python program (pycall) that then calls out to a C program (FCinterfaceCcode).  That's bad programming.  Streamline the program and get rid of python.

You also definitely need to fix whatever the issue is that requires running the program as root.  We can't have programs like this run as root.

 I tried making these changes but there was a problem with R pi boot again.I now know how to bypass the python code using IOC.I will make these changes once the problem with the Pi is fixed.

  10276   Sat Jul 26 13:38:34 2014 JamieUpdateGeneralData Acquisition from FC into EPICS Channels

Quote:

 I succeeded in creating a new channel access server hosted on domenica ( R Pi) for continuous data acquisition from the FC into  accessible channels. For this I have written a ctypes interface between EPICS and the C interface code to write data into the channels. The channels which I created are:

C1:ALS-X-BEAT-NOTE-FREQ

C1:ALS-Y-BEAT-NOTE-FREQ

 

The scripts I have written for this can be found in:

db script in:     /users/akhil/fcreadoutIoc/fcreadoutApp/Db/fcreadout.db

 Python code:  /users/akhil/fcreadoutIoc/pycall

C code:          /users/akhil/fcreadoutIoc/FCinterfaceCcode.c

I will give the standard channel names(similar to the names on the channel root)once the testing is completed and confirm that data from FC is consistent with the C code readout. Once ready I will run the code forever so that both the server and data acquisition are in process always.

Yesterday, when I set out to test the channel, I faced few serious issues in booting the raspberry pi. However, I have backed up the files on the Pi and will try to debug the issue very soon( I will test with Eric Q's R Pi).

To run these codes one must be root ( sudo python pycall, sudo ./FCinterfaceCcode)  because the HID- devices can be written to only by the root(should look into solving this issue). 

Instructions for Installation of EPICS, and how to create channel server on Pi will be described in detail in 40m Wiki ( FOLL page).

 

controls@rossa|~ 2> ls /users/akhil/fcreadoutIoc
ls: cannot access /users/akhil/fcreadoutIoc: No such file or directory
controls@rossa|~ 2> 

This code should be in the 40m SVN somewhere, not just stored on the RPi.

I'm still confused why python is in the mix here at all.  It doesn't make any sense at all that a C program (EPICS IOC) would be calling out to a python program (pycall) that then calls out to a C program (FCinterfaceCcode).  That's bad programming.  Streamline the program and get rid of python.

You also definitely need to fix whatever the issue is that requires running the program as root.  We can't have programs like this run as root.

  10275   Sat Jul 26 13:10:14 2014 AkhilUpdateGeneralThermal Actuator Transfer Functions

Koji said that the method we used for X-arm thermal actuator TF measurement was not correct and suggested us to make measurements separately for high and low frequencies( ensuring coherence at those frequencies is high).

(Edit by KA: The previous measurements for X/Y arm thermal actuators were done with each arm individually locked. This imposes the MC stability to the arm motion. The MC stability is worse than the arm stability due to shorter length and more number of the mirrors. Thus the arm motions were actually amplified rather than stabilized. The correct configuration was to stabilize MC using the other arm and control the measurement arm with the arm cavity length.)

So I and Eric Q took some improved TF measurements last night for the X-arm. The input excitation and the filters used were similar to that of the previous measurement . The attached are the TF plots showing two different frequency measurements.The data was saved and will be used to generate a complete TF. The attached (TFX_new.pdf)shows the independent TF measurement for X-arm temperature actuator. The black legend shows the TF at high frequencies(>1 Hz) and the red at low frequencies(<1 Hz). The final TF plots( from the data) will be posted in my next elog. 

We also made the measurements needed for calibration of these actuator Transfer functions. For this we gave some excitation for the arm length( separately for X arm and Y arm) and measured the PZT response. I will eLog with the details of the measurement and results shortly.

Attachment 1: TFX_new.pdf
TFX_new.pdf
  10274   Sat Jul 26 10:12:19 2014 AkhilUpdateGeneralData Acquisition from FC into EPICS Channels

 I succeeded in creating a new channel access server hosted on domenica ( R Pi) for continuous data acquisition from the FC into  accessible channels. For this I have written a ctypes interface between EPICS and the C interface code to write data into the channels. The channels which I created are:

C1:ALS-X-BEAT-NOTE-FREQ

C1:ALS-Y-BEAT-NOTE-FREQ

 

The scripts I have written for this can be found in:

db script in:     /users/akhil/fcreadoutIoc/fcreadoutApp/Db/fcreadout.db

 Python code:  /users/akhil/fcreadoutIoc/pycall

C code:          /users/akhil/fcreadoutIoc/FCinterfaceCcode.c

I will give the standard channel names(similar to the names on the channel root)once the testing is completed and confirm that data from FC is consistent with the C code readout. Once ready I will run the code forever so that both the server and data acquisition are in process always.

Yesterday, when I set out to test the channel, I faced few serious issues in booting the raspberry pi. However, I have backed up the files on the Pi and will try to debug the issue very soon( I will test with Eric Q's R Pi).

To run these codes one must be root ( sudo python pycall, sudo ./FCinterfaceCcode)  because the HID- devices can be written to only by the root(should look into solving this issue). 

Instructions for Installation of EPICS, and how to create channel server on Pi will be described in detail in 40m Wiki ( FOLL page).

 

  10273   Fri Jul 25 17:28:31 2014 HarryUpdateGeneralFOL Box and PER Update

 Purpose

We're putting together a box to go into the 1X2 rack, to facilitate the frequency counters, and Raspberry Pi that will be used in FOL.

Separately, I am working on characterizing the Polarization Extinction Ratio of the PM980 fibers, for further use in FOL.

What's Been Done

The frequency counters have been mounted on the face of the box, and nylon spacers installed in the bottom, which will insulate the RPi in the future, once it's finally installed.

FOLBox.png

In regard to the PER setup, there is an issue, in that the mounts which hold the collimators rotate, so as to align the axes of the fibers with the polarization of the incoming light.

This rotational degree of freedom, however, isn't "sticky" enough, and rotates under the influence of the stress in the fiber. (It's not much, but enough.)

This causes wild fluctuations in coupled power, making it impossible to make accurate measurements of PER.

What's Next

In the FOL box's case, we've ordered a longer power cable for the raspberry pi (the current one is ~9 inches long).

Once it arrives, we will install the RPi, and move the box into its place in the rack.

In the case of the PER measurement, we've ordered more collimator mounts//adapters, which will hopefully give better control over rotation.

 

  10272   Thu Jul 24 19:28:43 2014 AkhilUpdateGeneralThermal Actuator Transfer Functions

 As a part of temperature actuator characterization, today Eric Q and I made some measurements for the open loop TF of both the X-arm and Y-arm  thermal actuators. 

For this, we gave an input  of random excitation for the temperature offset input( since we faced some serious issues when we gave in Swept sine yesterday) and observed the PZT actuation signal keeping the arm to be locked all the time of our measurements and ensuring that the PZT signal doesn't saturate.

The  channels used for the measurement were  C1:ALS-X_SLOW_SERVO2_EXC as the input and C1:ALS-X_SLOW_SERVO1_IN1  as the output.

The random noise used for the measurement :

Y-ARM:  Gain- 6000;  Filter - butterworth-first order - band-pass filter with start frequency= 1 Hz stop frequency = 5 Hz.

X-ARM: Gain -3000; Filter - butterworth- first order- band-pass filter with start frequency 3 Hz and stop frequency = 30 Hz and  notch(1,10,20).

The Y-ARM measurement was stable but for the X-ARM, the PZT was saturating too often so Eriq Q went inside the lab and placed a 20dB attenuator in the path of the  X-ARM PZT signal readout to carry out the stable measurements.

The units of the TF of these measurements are not calibrated and are in count/count. I will have to calibrate the units by measuring the PZT count by changing the cavity length so that I can get a standard conversion into Hz/count. I will elog the calibrated TFs in my next elog after I take the cavity length and PZT TFs.

The attached are the bode plots for both the X-ARM and Y-ARM thermal actuators(non-calibrated). I will work on finding the poles and zeroes of this system once I finish calibration of the TF measurements.

Attachment 1: TF-X-ARM.pdf
TF-X-ARM.pdf
Attachment 2: TF-Y-ARM.pdf
TF-Y-ARM.pdf
  10271   Thu Jul 24 17:37:19 2014 HarryUpdateGeneralCoupling Improvements plus PER Measurement Setup

Quote:

 Purpose

We wanted to improve the coupling into the fibers, because it's very rarely good enough to take measurements with, as the beam is obscured by random noise.

Additionally, we want to add some things to the current setup in order to better measure Polarization Extinction Ratio.

What Was Done

After flailing for several hours, Koji helped me couple the NPRO light into the fiber, using the fiber illuminator for alignment. The coupled optical power immediately jumped from 0-1uW to 5.6mW (around 11% coupling).

Q and I discussed the setup for measuring PER. In addition to the current setup, we added a half wave plate to control the angle of the polarization, in addition to the existing quarter wave plate, which corrects the beam for ellipticity.

PERSetup.png

Once everything was coupled, I started minimizing S-Polarization coming out of the first polarizing beam splitter, and maximizing the P-Polarization entering the fibers.

I did this by first varying the Quarter Wave plate to eliminate as much S Polarization as possible, and then, maintaining a constant differential in angle between QWP and HWP, I rotated them both to maximize power coupled into the fibers.

I measured 0.2 mW of S-Polarization, and 54.3 mW of P-Polarization.

At this point, a locking effort started, and I had to leave the 40m.

Moving Forward

 

Tomorrow, I would like to finish the setup of the PER measurement design. That is to say, add a collimator to the other end of the fiber, and align it with the second PBS.

And, of course, take a measurement of the Polarization Extinction Ratio of the fiber.

To eventually be implemented in Frequency Offset Locking. 

 

Today, I encountered a problem with the stage that holds the coupler, in that its ability to rotate unchecked causes coupling to degrade over time due to torsion in the fibers. Our solution was to stress-relieve the fiber with a clamp.

Unfortunately, this also meant losing coupling completely. It was re-coupled at up 72% efficiency. (Subsequent changes in the setup have decreased that to ~24%)

When I took preliminary measurements of the PER, it was significant, which was unexpected. Upon further discussion with Q, we concluded that since the fiber's fast axis hadn't been aligned with the light's polarization, I was getting multiple polarizations out the end of the fiber.

Subsequent measurements of the power contained in the two polarizations of the output light gave about 0.8% S-Polarization introduced by the fiber.

Tomorrow

I would like to find another collimator holder, to hold the output side of the fiber.

Also, I will spend more time aligning the fiber axes, and the second PBS in order to get a better (read: more reasonable) measurement of PER.

  10270   Thu Jul 24 14:20:30 2014 SteveUpdateSUSPRM & other oplev gain settings checked

 The PRM sus gains checked OK

All other suspension oplev gains setting were checked out OK

 

Attachment 1: PRMgainsSensors.png
PRMgainsSensors.png
  10269   Thu Jul 24 13:01:39 2014 ericqUpdateSUSPRM OPLEV!

 Here's a fun fact: since the great computer failure of June2014, the PRM Oplev gains have been ZERO.

oops.png

arrrrggggh

I've restored the gains to their old values, and measured the loop TFs.

PRMOLPIT.pdfPRMOLYAW.pdf

 

 

 

 

  10268   Thu Jul 24 09:18:15 2014 SteveUpdateGeneralfour days
Attachment 1: 4days.png
4days.png
  10267   Wed Jul 23 23:43:28 2014 ericqUpdateLSCLocking efforts; Wrath of the Mode Cleaner

 [Koji, ericq]

We were working on getting back into the locking groove tonight.

The POP2F and REFL3F demod angles needed some tuning to lock the PRC reliably. The green alignments were mostly fine, the X end PZT ASS works reasonably well. Suspensions, especially the ITMs, seemed to be drifting a fair deal; today was fairly hot out, I guess.

We only got to the point of attempting the SqrtInv handoff once (which failed because I forgot to check the filter bank offsets). This was because the Mode Cleaner refused to stay locked longer than ~5-10 minutes at a time. We adjusted the MC and FSS servo offsets by the usual means, but this didn't make a difference.

We discussed and decided that the time is right to roll up our sleeves and dig into the MC loop, and try to figure out why these intermittent times of unreliability keep cropping up. We will check out the servo board, and see if we can find the missing phase than Evan observed, as well as characterize the FSS/PZT crossover, and investigate what kind of conditions we may create that cause the PC to saturate.

  10266   Wed Jul 23 19:30:34 2014 NichinUpdateElectronicsTime delay in the RF multiplexer (Rack 1Y1)

A time delay can be modeled as the exponential transfer function :  e(-sTd)  as seen HERE . Therefore the slope of the phase gives us the time delay.

The transfer function of RF multiplexer in rack 1Y1 (NI PXI-2547) was fit to an ideal delay function e(-sTd) , with Td = 59 ns.

The plots shows the actual data, fit data and data after correction using the ideal model stated above.

Conclusion:

Delay the RF Multiplexer is approximately 59 ns. This value can be used to correct the phase in measurements of transimpedance for each PD by dividing out the ideal transfer function for time delay.

 

Attachment 1: RFmux1.pdf
RFmux1.pdf
Attachment 2: RFmux2.pdf
RFmux2.pdf
  10265   Wed Jul 23 18:53:11 2014 NichinUpdateElectronicsTime delay in RG405 coaxial cables

 A time delay can be modeled as the exponential transfer function :  e(-sTd)  as seen HERE . Therefore the slope of the phase gives us the time delay.

A RG405 coaxial cable, exactly 5.5 meters in length, was fit to an ideal delay function e(-sTd) , with Td = 150 ns.

The plots shows the actual data, fit data and data after correction using the ideal model stated above.

Conclusion:

Delay in RG405 cables is approximately 27.27 ns per meter. This value can be used to correct the phase in measurements of transimpedance for each PD by dividing out the ideal transfer function for time delay.

[EDIT: This looks like we have about 12 % the speed of light inside the RF cables. Too small to be true. I will check tomorrow if the Network analyzer itself has some delay and update this value.]

The varying attenuation of about 1dB due to the cable is not compensated by this. We need to separately include this.

Things to do:

1) Get the length of RF cables that is being used by each PD, so that the compensation can be made.

2) Calculate the attenuation and delay caused by RF multiplexer and Demodulator boards. Include these in the correction factor for transimpedance measurements. 

 

 

 

 

 

 

 

 

 

 

Attachment 1: RFcable1.pdf
RFcable1.pdf
Attachment 2: RFcable2.pdf
RFcable2.pdf
  10264   Wed Jul 23 17:54:51 2014 HarryUpdateGeneralCoupling Improvements plus PER Measurement Setup

 Purpose

We wanted to improve the coupling into the fibers, because it's very rarely good enough to take measurements with, as the beam is obscured by random noise.

Additionally, we want to add some things to the current setup in order to better measure Polarization Extinction Ratio.

What Was Done

After flailing for several hours, Koji helped me couple the NPRO light into the fiber, using the fiber illuminator for alignment. The coupled optical power immediately jumped from 0-1uW to 5.6mW (around 11% coupling).

Q and I discussed the setup for measuring PER. In addition to the current setup, we added a half wave plate to control the angle of the polarization, in addition to the existing quarter wave plate, which corrects the beam for ellipticity.

PERSetup.png

Once everything was coupled, I started minimizing S-Polarization coming out of the first polarizing beam splitter, and maximizing the P-Polarization entering the fibers.

I did this by first varying the Quarter Wave plate to eliminate as much S Polarization as possible, and then, maintaining a constant differential in angle between QWP and HWP, I rotated them both to maximize power coupled into the fibers.

I measured 0.2 mW of S-Polarization, and 54.3 mW of P-Polarization.

At this point, a locking effort started, and I had to leave the 40m.

Moving Forward

 

Tomorrow, I would like to finish the setup of the PER measurement design. That is to say, add a collimator to the other end of the fiber, and align it with the second PBS.

And, of course, take a measurement of the Polarization Extinction Ratio of the fiber.

To eventually be implemented in Frequency Offset Locking.

  10263   Wed Jul 23 11:54:27 2014 NichinUpdateElectronicsCharacterization of demodulator boards.

Quote:

 

I repeated the exact steps above and made sure everything was back where it should be after I was done.

Reason I had to retake the measurements:

My script for acquiring data from the AG4395A network analyzer was such that it first acquired the magnitude data from channel 1 and then recorded phase data from channel 2 without holding its trace. Hence the phase and magnitude data were not exactly in sync with each other. So, when I tried to fit the data to a model using vector fitting, I ended up with very bad results.

I have now changed every single script relating to the network analyzer to just get the real and imaginary data in one go and then calculate the phase using this data.

The fitting process is now in progress and results will be up shortly.

The plots in the previous Elog includes delay and a little attenuation by RF cables and the RF mux.

Today I separately calculated the delay and attenuation for an RG405 cable (550 cm) and the RF mux(using really small RF cables). These delays should be accounted for when fitting the transfer function of Demodulator boards and transimpedance of PDs.

The plots are in both semilogx and linear.

Attachment 1: 1.pdf
1.pdf 1.pdf
Attachment 2: 2.pdf
2.pdf 2.pdf
  10262   Wed Jul 23 11:32:04 2014 KojiUpdateLSCIFO warming up

Alone with the IFO. Started from some conversation with it.

Some ALS trials: Found the Y-end green alignment was terrible. In fact the end green set up is terrible.
Unfixed optics, clipping/fringing in the faraday, unstable suprema mounts which is unnecessarily big.

Eventualy I stopped touching the end alignment. Run ALS to see the stability of the things.
This is a performance confirmation after some touching of the ALS electronics by Manasa/SURFs

The sensing noise levels of the ALSs looks the same as before.

The intensity noise of the transmission was also checked. They are not RIN but very close to RIN
as the DC was the unity for both arms.

The X arm has worse ALS noise level and RIN.
Although I forgot to turn off the HEPA flow at the south bench during the measurement. Gurrr.

Attachment 1: 140722_ALS.pdf
140722_ALS.pdf
  10261   Wed Jul 23 11:15:54 2014 AkhilUpdateElectronicsInstallation of FCs in the 40m

 As a part of installation of two(X-ARM and Y-ARM) frequency counters in the 40m, I have tested their performance when using them both on a single Raspberry Pi. The timing plots are attached. There are almost no timing issues in this configuration and it can be said that there is no harm using both of the FCs on the same platform.

We will be installing the FC box inside the lab and carry out few tests with RF mon beat note inputs.

Attachment 1: Timingwith2FCs.png
Timingwith2FCs.png
  10260   Wed Jul 23 10:40:23 2014 NichinUpdateGeneralWeekly Update

To do:

  1. Measure and calibrate out  attenuation and phase changes due to RF cables in the PDFR system.
  2. Create a database of canonical plots for comparison each time new data is acquired.
  3. Vector fitting or LISO fitting of transimpedance curves.

Does not require time from a lab expert.

  10259   Wed Jul 23 10:39:18 2014 SteveUpdateCamerasvideo quad processors replaced

Quad processor 2 & 3 were replaced.

  10258   Wed Jul 23 02:01:15 2014 JenneUpdateLSCRIN in arm transmission - revised calibration

 

As Koji pointed out, I messed up the calibration.  However, fixing it doesn't change things that much.

From this calibration by Yuta, the Xarm ALS calibration is 54 deg / MHz, or 19.17 kHz / deg.  So, I multiply my data which is in these degree units by 19.17e3 to get Hz.  Then I use delta_f / f = delta_L / L to convert to meters.  f = c / lambda_green, and L = 37.5 meters. 

This only changes the calibration by about 10-15%.  It still looks like the ALS noise is well above the RIN level of the sqrtInv signal.

TransSqrtInvRIN_vs_ALSsensing_18July2014.png

  10257   Tue Jul 22 23:10:12 2014 AkhilUpdateGeneralWeekly Update

 Work Done:

  • Created a Channel Access Server on the Raspberry Pi  to write data from the FC into EPICS Channel.
  • Completed characterization and noise estimation of the FC counter with improved timing.
  • Started installation of FC inside the 40m.

Plans for this Week:

  • Testing how well the FC can replace the spectrum analyzer which is in the control room. For this I have asked Steve to order  an RF adder/combiner to see how frequency counter responds to two RF signals at different frequencies(much like the RF signal fed to the spectrum analyzer) .
  • Complete the installation of FC insode the 40m and start initial testing.
  • Characterization of the Temperature Actuator and initial PID loop design.

Inside the 40m Lab:

  • I will have to go inside the 40m lab this week for routing the RF mon cables to the FC box(in detail:http://nodus.ligo.caltech.edu:8080/40m/10163) .
  • Also to setup for characterization of the temperature actuator, I will be required to go inside the lab in this week.
  10256   Tue Jul 22 17:45:11 2014 HarryUpdateGeneralWeekly Update

 The Past Week

 

I spent the past week coupling NPRO light into the fibers, and subsequently measuring the fiber mode profile using the beam profiler.

The Next Week

In the next week, I plan to at least do measurements of the Polarization Extinction Ratio of the fibers.

Materials

My current optical setup, plus an additional polarizing beam splitter (have it).

  10255   Tue Jul 22 16:26:04 2014 HarryUpdateGeneralFiber Mode Measurement

I repeated this process once more, this time using the computer controlled stage that the beam profiler is designed to be mounted to.

These data//fitting appears to be within error bars. The range of my measurements was limited when the beam width was near the effective aperture of the profiler.

This latest trial yielded a waist of 4um, located 2.9 mm inside the fiber for the X profile, and 3.0mm inside the fiber for the Y profile.

fiberModeProfile3.png

Code is attached in fiberModeMeasurement4.zip. Note that the z=0 point is defined as the end of the fiber.

Attachment 2: fiberModeMeasurement4.zip
  10253   Tue Jul 22 15:54:19 2014 ericqUpdateSUSITMY Oplev Recentered

 ITMY oplev was nearly clipping in yaw, causing wonky behavior (POY lock popping in and out frequently). I recentered it and the arm is locking fine now. 

  10252   Tue Jul 22 15:50:35 2014 NichinSummaryElectronicsCharacterization of demodulator boards.

Quote:

Rack 1Y2, I took transfer function measurements for each of the following demodulator boards: REFL11, REFL33, REFL55, REFL165, AS55, POP22, POX11 and POY11.

What I did:

1) Removed the wire at PD Input to demodulator board.

2) Put the MOD output from network analyzer into PD input of board.

3) Ran a sweep from 100kHz to 100MHz.

4) Measured the transfer function between PD RF MON and PD Input. (The PD RF MON signal came out of the RF multiplexer, so the mux is included as well )

5) Put the original wire back at PD Input.

Results:

The plots clearly show an attenuation of 20dB (factor of 10) for all the demodulator boards. This explains why my transimpedance measurements are off by 10 times.

Note: for REFL 165, there was an extra 100MHz high pass filter installed at PD Input. I did not remove this and made my measurements along with this.

To Do:

a) Modify the PDFR system to calibrate out this attenuation.

b) Measure the transfer function between the input and output of RF mux, so that we can have just the transfer function between PD input an PD RF MON (for documentation's sake)

 

I repeated the exact steps above and made sure everything was back where it should be after I was done.

Reason I had to retake the measurements:

My script for acquiring data from the AG4395A network analyzer was such that it first acquired the magnitude data from channel 1 and then recorded phase data from channel 2 without holding its trace. Hence the phase and magnitude data were not exactly in sync with each other. So, when I tried to fit the data to a model using vector fitting, I ended up with very bad results.

I have now changed every single script relating to the network analyzer to just get the real and imaginary data in one go and then calculate the phase using this data.

The fitting process is now in progress and results will be up shortly.

Attachment 1: Demodulators_TF.pdf
Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf
  10251   Tue Jul 22 08:36:08 2014 EvanUpdateIOOMC servo TFs

Quote:

[Rana, Evan]

This morning we took several TFs of the MC servo board using the HP4395A.

The 4395 source was teed, with one output of the tee going to 4395 R and the other output going to the board's IN1. We then took TFs of (4395 A) / (4395 R), where 4395 A was one of the following four points on the servo board:

  • OUT2
  • A TEST1
  • B TEST1
  • SERVO

For each of these points, we took a TF at two gain settings: IN1 and VCO gains both at 0 dB, and then IN1 and VCO gains both at 20 dB.

Before doing these measurements, we calibrated out the cable delay. Additionally, SERVO was always loaded with 50 Ω—either from the 4395 or from a terminator.

The attached png shows the servo board settings when these TFs were taken with the 0 dB gain settings. The settings for the 20 dB measurements are identical, except for the higher IN1 and VCO gains.

Using the modified schematic (40m:10250), I've made a plot of the TFs I expect for GIN1 = GVCO = 0 dB, taking into account our 50 Ω loading of the board.

Evidently I'm somehow missing a factor of 2 in the gain of the overall TF, but the shapes of the expected vs. measured magnitudes agree quite well.

At 1 MHz, I expect we should have accumulated about 80 degrees of phase going through the servo board. In reality, we appear to have lost more like 105 degrees.

Attachment 1: MCtfExpectations.pdf
MCtfExpectations.pdf
  10250   Tue Jul 22 08:24:42 2014 EvanUpdateElectronicsMC servo card: modified schematic

Quote:

Patient One for this new system will be the MC servo card. The DCC number is D1400242. Currently, v1 is just the original drawing with no modifications. I've updated the DCC document tree at E1400326 accordingly.

It looks like we can use Jenne's information in 40m:9892 to deduce the modifications that have been made (alternatively, someone can just pull the board and examine it on the bench).

The attached zip file has a modified schematic of the MC servo card (011/MC), as deduced from Jenne's photos. Someone should go through and verify that the schematic is correct. Then it can go on the DCC as D1400242-v2.

To modify the schematic, I used Inkscape (the svg files for each sheet are included in the zip file). Then to generate the pdf, I ran

for i in sheet*.svg; do inkscape -A "${i/svg/pdf}" "$i"; done

pdftk sheet*.pdf cat output D1400242

Attachment 1: D1400242.zip
  10249   Mon Jul 21 18:08:19 2014 HarryUpdateGeneralFiber Mode Measurement

Quote:

 Purpose

The idea was to measure the profile of the light coming out of the fiber, so we could have knowledge of it for further design of measurement apparatuses, for characterization of the fibers' properties.

Methods

The method was the same as the last time I tried to measure the fiber mode.

This time I moved the beam profiler in a wider range along the z-axis.

Additionally, I adjusted the coupling until it gave ~1mW through the fiber, so the signal was high enough to be reliably detectable.

Measurements were taken in both X and Y transections of the beam.

The range of movement was limited by the aperture of the beam profiler, which cuts off at 9mm. My measurements stop at 8.3mm, as the next possible measurement was beyond the beam profiler's range.

fiberModeMeasurement.png

Analysis

I entered my data into A La Mode, which gave me a waist of 5um, at a location of z = -0.0071 m, that is to say, 7.1mm inside the fiber.

Note that in the plot, data points and fits overlap, and so are sometimes hard to distinguish from each other.

Code is attached.

fiberModeFit2.png

Moving Forward

Using this data, I will begin designing setups to measure fiber characteristics, the first of which being Polarization Extinction Ratio.

Eventually, the data collected from these measurements will be put to use in the frequency offset locking setup.

 Edit

 

 

The previous data were flawed, in that they were taken in groups of three, as I had to move the micrometer stage which held the beamscan between holes in the optical table.

In order to correct for this, I clamped a straightedge (ruler) to the table, so I could more consistently align the profiler with the beam axis.

These data gave a waist w_o = 4um, located 6mm inside the fiber. While these figures are very close to what I would expect (3.3um at the end of the fiber) the fitting still isn't as good as I would like.

The fit given by ALM is below.

fiberModeMeasurement3.png

Moving Forward

I would like to get a stage//rail so I can align the axes of the beam and profiler more consistently.

I would also like to use an aperture the more precisely align the profiler aperture with the beam axis.

Once these measurements have been made, I can begin assembling the setup to measure the Polarization Extinction Ratio of the fiber.

  10248   Mon Jul 21 17:32:43 2014 ericqSummaryLSCArm losses

Quote:

From the last plot:

- Subtracting the offset of 0.0095, the modulation depth were estimated to be 0.20 for 11MHz, 0.25 for 55MHz

- Carrier TEM00 1.0, 1st order 0.01, 2nd order 0.05, 3rd order 0.002, 4th order 0.004

==> mode matching ~93%, dominat higher order is the 2nd order (5%).

Eric: now we have the number for the mode matching. How much did the cavity round-trip loss be using this number?

Using these numbers for both arms (Modulation takes away .2*.25 = 5% power, mode matching takes away 7% after that), I get the following from my data from March:

Xarm loss is 561.19 +/- 14.57 ppm

Yarm loss is 130.67 +/- 18.97 ppm

Obviously, the Xarm number looks very fishy, but its behavior was qualitatively very different when I took the data. ASDC would change from ~0.298 to ~0.306 when the Yarm was locked vs. misaligned, whereas the xarm numbers were .240 to .275. 

In any case, I'll do the measurement again tomorrow, being careful with offsets and alignment; it won't take too long. 

  10247   Mon Jul 21 13:58:33 2014 ericqUpdateIOOMC autolocker acting up

The autolocker claimed it was running and blinking, but not doing anything (i.e. lock bit was not updating and no switches or sliders being touched)

After stopping and starting it a number of times, it began working again, through no real changes of my own. I'm a little mystified as to what the problem was... keep an eye out.

  10246   Mon Jul 21 12:16:27 2014 AkhilSummaryElectronicsFilters used inside the Frequency Counter

The expected bode plots for such a filter with L= 4 is attached and compared with the measured.

RXA: When comparing two things, please put them onto the same plot so that they can be compared.

Attachment 1: FC_TF_Characterization.png
FC_TF_Characterization.png
  10245   Mon Jul 21 10:51:06 2014 SteveUpdateVACN2 supply run out

Interlock closed valve V1, V4, V5 and VM1 when the nitrogen supply run out. The IFO pressure rose to P1 1 mTorr

In order to recover Vacuum Normal valve configuration I did the following:

Replaced both nitrogen cylinders. Confirmed pneumatic nitrogen pressure 70 PSI.   Opened valves V4 and V5

At P2 < 1 mTorr, Maglev rotation 560 Hz , V1 was opened.

VM1 was opened when CC1 pressure dropped below < 1e-5 torr

 

Please  take a look at the N2 cylinders pressure on Friday to insure that there is enough for the week end.

The daily consumption is 600-700 PSI

Attachment 1: outofN2.png
outofN2.png
  10244   Mon Jul 21 10:30:38 2014 HarryUpdateGeneralFiber Mode Measurement

 Purpose

The idea was to measure the profile of the light coming out of the fiber, so we could have knowledge of it for further design of measurement apparatuses, for characterization of the fibers' properties.

Methods

The method was the same as the last time I tried to measure the fiber mode.

This time I moved the beam profiler in a wider range along the z-axis.

Additionally, I adjusted the coupling until it gave ~1mW through the fiber, so the signal was high enough to be reliably detectable.

Measurements were taken in both X and Y transections of the beam.

The range of movement was limited by the aperture of the beam profiler, which cuts off at 9mm. My measurements stop at 8.3mm, as the next possible measurement was beyond the beam profiler's range.

fiberModeMeasurement.png

Analysis

I entered my data into A La Mode, which gave me a waist of 5um, at a location of z = -0.0071 m, that is to say, 7.1mm inside the fiber.

Note that in the plot, data points and fits overlap, and so are sometimes hard to distinguish from each other.

Code is attached.

fiberModeFit2.png

Moving Forward

Using this data, I will begin designing setups to measure fiber characteristics, the first of which being Polarization Extinction Ratio.

Eventually, the data collected from these measurements will be put to use in the frequency offset locking setup.

Attachment 3: fiberModeMeasurement2.zip
  10243   Sun Jul 20 09:26:27 2014 EvanUpdateElectronicsMC servo card modifications in DCC

Quote:

[Rana, Jenne]

We have decided to keep better track (using new-fangled digital "computers") of our modifications to electronics boards. 

The idea will be to create a new DCC document for every electronics board (when we pull a board and modify it, it should receive this treatment) that we have, and that document will become a history of the board's life.  Version 1 will be a copy of the original drawing.  Version 2 should be a modified version of that drawing with the current situation.  All future versions should be modified from the most recent version, to reflect any changes.  Notes for each updated version should include an elog reference to the work, so that we know why we did things, and have a place to find photos of the actual modifications.  Elogs should also include a link to the DCC version.  DCC titles should include the phrase "40m Revisions" for ease of searching.

Patient Zero for this new system will be the PMC servo card.  The DCC number is D1400221.  As of this moment, this just has the V1 original drawing with no modifications.

This has been included in the 40m's DCC document tree that Jamie started back in November 2012.

Patient One for this new system will be the MC servo card. The DCC number is D1400242. Currently, v1 is just the original drawing with no modifications. I've updated the DCC document tree at E1400326 accordingly.

It looks like we can use Jenne's information in 40m:9892 to deduce the modifications that have been made (alternatively, someone can just pull the board and examine it on the bench).

  10242   Sat Jul 19 20:51:51 2014 KojiUpdateLSCRIN in arm transmission

Your calibration of the ALS signal should be revised.

The phase for the ALS is not an optical phase of the green but the phase of the phase tracker servo output.

The calibration of the phase tracker depends on the cable length of the delay line in the beat box.
It seems that we are based on this calibration. Which gives up ~19kHz/deg.

Or, equivalently, use C1:.....PHASE_OUT_HZ instead.

  10241   Sat Jul 19 17:36:44 2014 JenneUpdateLSCRIN in arm transmission

I looked at what the RIN contribution of the sqrtInv sensor is by locking the arms individually on IR using POX and POY.  I then took spectra of the sqrtInv channels.  For the Xarm, I had forced the triggering so that the QPD was being used as the transmission PD, while the Yarm was using the regular Thorlabs PD.  I also had the green lasers locked to the arms, and took beatnote spectra to see what the sensing noise of the beatnotes is, all at the same time.

For the sqrtInv channels, I used the Optickle calibration from elog 10187. For today's plot, I am using the calibration at about 1nm, since that is about where we are when we transition to the sqrtInv Thorlabs signal usually.

For the ALS channel, I was using the _FINE_PHASE_OUT signal, which is in units of degrees of phase for a single green wavelength.  So, since k * x = phi, I want the phase data to be converted to radians (2*pi/360), and use k = 2*pi / lambda_green.  So, doing some algebra, this gives me x = phi_degrees * lambda / 360 for my calibration. 

What I see in the plot is that the ALS sensing noise is pretty bad compared to the sqrtInv channels, so maybe we don't have to work so hard on the ISS this next week.  Also, the Thorlabs PD is much better than the QPDs, which maybe isn't so surprising since we have them set so that they have good SNR at higher power.

Anyhow, here's the plot:

TransSqrtInvRIN_vs_ALSsensing_18July2014.png

Also, here is the Thorlabs PD only, with single arm locked on RF, with the noise calibrated to different CARM offsets:

ThorlabsRIN_16July2014.png

  10240   Sat Jul 19 01:59:34 2014 HarryUpdateGeneralFiber Mode Measurement

Purpose 

We wanted to measure the mode coming out of the fibers, so we can later couple it to experimental setups for measuring different noise sources within the fiber. i.e. Polarization Extinction Ratio, Frequency Noise, Temperature Effects.

Methods

I used the beamscan mounted on a micrometer stage in order to measure the spot sizes of the fiber coupled light at different points along the optical axis, in much the same way as in the razorblade setup I used earlier in the summer.

fiberModeMeasurement.png

Analysis

I entered my data (z coordinates, spot size in x, spot size in y) into a la mode to obtain the beam  profile (waist size, location)

 fiberModeMeasurement1.png 

Code is attached in .zip file.

Moving Forward

After I took these measurements, Manasa pointed out that I need points over a longer distance. (These were taken over the range of the micrometer stage, which is 0.5 inches.)

I will be coming in to the 40m early on Monday to make these measurements, since precious beamscan time is so elusive.

Eventually, we will use this measurement to design optical setups to characterize Polarization Extinction Ratio, Frequency Noise, and temperature effects of the fibers, for further use in FOL.

Attachment 3: fiberModeMeasurement1.zip
  10239   Fri Jul 18 19:32:50 2014 AkhilSummaryElectronicsFilters used inside the Frequency Counter

 

 Thanks Koji , for your  hint for the brain teasing puzzle. I was looking into Filters that are usually used in devices like counters, DSO and other scopes. I found that , to improve the quality of the measurement one of the best approach  is averaging. I looked deeper into averaging and found out this:

There are two general use-cases for averaging . The first, successive sample averaging, takes a single acquisition and averages between its samples. The second, successive capture averaging, combines the corresponding  samples of multiple captures to create a single capture. Successive sample averaging is also called boxcar filtering or moving average filtering. In an implementation of this type of averaging each output sample represents the average value of M consecutive input samples. This type of averaging removes noise (one of the reasons the noise level was not bad: http://nodus.ligo.caltech.edu:8080/40m/10151) by decreasing the device's bandwidth(could be one of the reasons why the FC operates in 4 different frequency ranges). It applies an LPF function with a 3dB point approximated by  0.433 * s / M, where M is the number of samples to be averaged, and s is the sample rate in samples per second. 

Now I tried verifying the 3 dB points in the gain plots I generated :

For 1 s Sampling time : the 3 dB point for such a Boxcar filter should be at 0.433* 1/M. If we assume that it averages for 2 samples, M=2 which gives the 3dB point at 0.288 Hz but occurs somewhere between 0.3 and 0.4 Hz.  (http://nodus.ligo.caltech.edu:8080/40m/140619_120548/GainVsFreq.png)

For 0.1s Sampling time: the 3dB point should be at 2.17 Hz and in reality is 2.5 Hz(http://nodus.ligo.caltech.edu:8080/40m/140701_211904/gain.png).

Also, This type of filter will have very sharp nulls at frequencies corresponding to signals whose periods are integer sub-multiples of M/s. As seen my previous plots (http://nodus.ligo.caltech.edu:8080/40m/10118 , http://nodus.ligo.caltech.edu:8080/40m/10070) there are sharp nulls at frequencies

0.4 Hz for 1S sampling time and

at 1.5 Hz,3 Hz for 0.1 S sampling time as correctly predicted.

The moving average filter is  L-sample moving average FIR, with the frequency response as:   H(ω) = (1/L) (1 − e− jω L)/(1 − e− jω)..

There is an overall delay of (M - 1)/2 samples from such a length-M causal FIR filter. 

The expected bode plots for such a filter with L= 5 is attached(attachment 2).

Attachment 1: TheoreticalGainPlot.png
TheoreticalGainPlot.png
Attachment 2: TFexpected.png
TFexpected.png
  10238   Fri Jul 18 17:10:57 2014 NichinSummaryElectronicsCharacterization of demodulator boards.

Rack 1Y2, I took transfer function measurements for each of the following demodulator boards: REFL11, REFL33, REFL55, REFL165, AS55, POP22, POX11 and POY11.

What I did:

1) Removed the wire at PD Input to demodulator board.

2) Put the MOD output from network analyzer into PD input of board.

3) Ran a sweep from 100kHz to 100MHz.

4) Measured the transfer function between PD RF MON and PD Input. (The PD RF MON signal came out of the RF multiplexer, so the mux is included as well )

5) Put the original wire back at PD Input.

Results:

The plots clearly show an attenuation of 20dB (factor of 10) for all the demodulator boards. This explains why my transimpedance measurements are off by 10 times.

Note: for REFL 165, there was an extra 100MHz high pass filter installed at PD Input. I did not remove this and made my measurements along with this.

To Do:

a) Modify the PDFR system to calibrate out this attenuation.

b) Measure the transfer function between the input and output of RF mux, so that we can have just the transfer function between PD input an PD RF MON (for documentation's sake)

 

Attachment 1: Demodulators_TF.pdf
Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf Demodulators_TF.pdf
  10237   Fri Jul 18 16:52:56 2014 AndresUpdate40m Xend Table upgradeFInish Calculation on Current X-arm mode Matching

Quote:

Data and Calculation for the Xarm Current Mode Matching

Two days ago, Nick, Jenne, and I took a measurement for the Green Transmission for the X-arm. I took the data and I analyzed it. The first figure attached below is the raw data plotted. I used the function findpeaks in Matlab, and I found all the peaks. Then, by taking close look at the plot, I chose two peaks as shown in the second figure attached below. I took the ratio of the TEM00 and the High order mode, and I average them. This gave me a Mode Matching of 0.9215, which this value is pretty close to the value that I predicted by using a la Mode in http://nodus.ligo.caltech.edu:8080/40m/10191, which is 0.9343. Nick and I measured the reflected power when the cavity is unlocked and when the cavity is locked, so we measured the PreflUnLocked=52+1µW and PreflOnLocked=16+2µW and the backgroundNoise=0.761µW. Using this information we calculated  Prefl/Pin=0.297. Now, since Prefl/Pin=|Eref/Ein|2, we looked at the electric fields component by using the reflectivity of the mirror we calculated 0.67. The number doesn't agree, but this is because we didn't take into account the losses when making this calculation. I'm working in the calculation that will include the losses.

Today, Nick and I ordered the lenses and the mirrors. I'm working in putting together a representation of how much improvement the new design will give us in comparison to the current setup.

We want to be able to graphically see how much better it is the new optical table setup in comparison to the current optical table setup. In other words, we want to be able to see how displacement of the beam and how much angle change can be obtained at the ETM from changing the mirrors angles independently. Depending on the spread of the mirrors' vectors we can observe whether the Gouy phase is good. In the plot below, the dotted lines correspond to the current set up, and we can see that the lines are not spread from each other, which essentially mean that changing the angles of the two mirrors just contribute to small change in angle and in the displacement of the beam at the ETM, and therefore the Gouy phase is not good. Now on the other hand. The other solid lines correspond to the new setup mirrors. We can observe that the spread of the line of mirror 1 and mirror 4 is almost 90 degrees, which just implies that there is a good Gouy phase different between these two mirrors. For the angles chosen in the plot, I looked at how much the PZT yaw the mirrors from the elog http://nodus.ligo.caltech.edu:8080/40m/8912. In this elog, they give a plot in mrad/v for the pitch and yaw, so I took the range that the PZT can yaw the mirrors, and I converted into mdegrees/v and then I plotted as shown below. I plot for the current setup and for the new setup in the same plot. The matlab code is also attached below.

Attachment 1: OldAndNewSetupPlotsOfDisplacementAndAngleAtTheETM.png
OldAndNewSetupPlotsOfDisplacementAndAngleAtTheETM.png
Attachment 2: OldSetUpDisplacementAndNewSetup.m.zip
  10236   Fri Jul 18 15:21:12 2014 ericqUpdateComputer Scripts / ProgramsLocal Chiara backups

Quote:

I've also written a backup script at scripts/backup/rsync_chiara.backup which keeps its books in scripts/backup/rsync_chiara.backup.log 

I'm adding a entry to the root crontab on chiara to execute the script every day at 7am. 

 I had some syntax errors in the script that prevented the script from doing the right thing. The backup is now up to date, and the cronjob should work. 

  10235   Fri Jul 18 14:59:07 2014 EvanUpdateIOOMC servo TFs

[Rana, Evan]

This morning we took several TFs of the MC servo board using the HP4395A.

The 4395 source was teed, with one output of the tee going to 4395 R and the other output going to the board's IN1. We then took TFs of (4395 A) / (4395 R), where 4395 A was one of the following four points on the servo board:

  • OUT2
  • A TEST1
  • B TEST1
  • SERVO

For each of these points, we took a TF at two gain settings: IN1 and VCO gains both at 0 dB, and then IN1 and VCO gains both at 20 dB.

Before doing these measurements, we calibrated out the cable delay. Additionally, SERVO was always loaded with 50 Ω—either from the 4395 or from a terminator.

The attached png shows the servo board settings when these TFs were taken with the 0 dB gain settings. The settings for the 20 dB measurements are identical, except for the higher IN1 and VCO gains.

Attachment 1: mcServoTFSettings.png
mcServoTFSettings.png
Attachment 2: MCtfs.pdf
MCtfs.pdf
  10234   Thu Jul 17 22:08:14 2014 KojiUpdateGeneral1X2 Rack Changes

It sounds like the work was done carefully. Even so, Jenne or Manasa have to run the ALS (X and Y) to check if they are still functional.

ELOG V3.1.3-